OpenStudy (anonymous):

Can anyone please add this rational expression, and demonstrate and explain how to do so? Thanks. 2x-2/4x + 3x+3/2x

OpenStudy (freethinker):

I wouldn't ask the difference between the denominators that would be 4x - 2x = 2x I don't think this is the right approach just like when you are adding \[\frac{ 1 }{ 2 }+ \frac{ 1 }{ 4 }\] you don't ask what is the difference between the denominators DIFFERENCE in mathematical terms means subtraction, so it would be a very misleading term to use when adding fractions

OpenStudy (anonymous):

Right. But Im still confused as to how to solve it though